A sewage treatment plant has two inlet pipes to its settling pond. One pipe can fill the pond in 10 hours, the other pipe in 12 hours. If the first pipe is open for 5 hours and then the second pipe is opened, how long will it take for both pipes working together to fill the pond?
Answer by jorel1380(3719)

t/10+t/12=1/2
6t/60+5t/60=30/60
6t+5t=30
11t=30
t=30/11 more hours for both pipes to fill the pond..